# An Information Theoretic Perspective on Conformal Prediction

### v2

@@ 7 Experiments
In this section, we empirically study two applications of our theoretical results, namely conformal prediction with side information and conformal training with our upper bounds on the conditional entropy as optimization objectives. We focus our experiments on classification tasks since this is the most common setting in previous works in conformal training [8, 59].
@@ 7.1 Conformal Training
In Table 1, we report the empirical inefficiency on test data considering two SCP methods, threshold CP with probabilities (or THR) [54] and APS [53]—see Appendix G.1.2 for results with RAPS [4]. In all cases, our upper bounds proved effective loss functions to train efficient classifiers end-to-end and from scratch. For the simpler data sets (MNIST, Fashion-MNIST and EMNIST), all conformal training methods achieved similar results, but both ConfTr methods proved less consistent. This is noticeable in the oftentimes sharp difference in performance between THR and APS, since even after fine-tuning the hyperparameters (see Appendix G) some of the models fail to converge properly. For the remaining and more challenging data sets, both ConfTr variants lagged behind, probably because they do not provide a strong enough signal to train ResNets from scratch (on CIFAR data sets, Stutz et al. [59] only used ConfTr to fine-tune pretrained models). A similar observation applies to the Fano bound, where the only signal to the classifier comes through the relaxed prediction set size in (6). Finally, we also observe that our conformal training objectives outperform the baselines even if $\alpha$ values at test time differ from the one used for training, as we show in Table 10 in Appendix G.1.1.
@@ 7.3 Federated Learning (FL)
One marked difference between Tables 1 and 3 is that the simple Fano bound (6), which lagged behind the DPI bound and its model-based counterpart in the centralized setting, achieved the best results on the federated setting. We hypothesize this could be due to overfitting of the local optimization procedures to the individual data distribution of each device, which hurts the convergence of the global model. This effect is exacerbated on CIFAR100, where we have 500 devices, each of which with very few data points. The simple Fano bound is less vulnerable to such overfitting since it relies on the main classifier to a much lesser degree. Finally, in almost all cases, our bounds outperformed the CE loss, reassuring the potential of conformal training. Moreover, the inclusion of side information reduced inefficiency in all settings, and markedly so in a few instances. This confirms the effectiveness of our side information approach in a complex and practically relevant scenario, like federated learning.
@@ 8 Conclusion
In this work, we established a link between notions of uncertainty coming from conformal prediction and information theory (or more precisely variable-size list decoding). We proved that one can use split conformal prediction methods to upper bound the conditional entropy of the target variable given the inputs, and that these upper bounds form principled objectives for conformal training. We empirically validated our approach to conformal training, with strong results in both centralized and federated settings. Furthermore, the information-theoretic perspective also offers a simple yet rigorous approach to incorporate side information into conformal prediction, which we experimentally show leads to better predictive efficiency. To the best of our knowledge, this is the first attempt at connecting information theory and conformal prediction. Given the limited communication between these two research communities thus far, we expect our work to incite a fruitful exchange of not only ideas but also theory and algorithms between these two research domains. We see an extension of our methods to the regression setting, as a particularly promising avenue for future work.
@@ Appendix A Broader Impact, Limitations and Other Remarks
This work explores the connection between conformal prediction and information theory, with the end goal of advancing the current state of the art of uncertainty estimates in machine learning. With that in mind, we believe its potential societal consequences are chiefly positive, since our work might contribute to a larger adoption of uncertainty estimates, in a number of safety-critical applications. Notwithstanding, conformal prediction, like any other uncertainty estimation method, should be applied with care and a proper understanding of the provided guarantees so as not to create an illusion of safety when there is none. That is why our work aims to develop new methods and algorithms from first principles, with the hope of providing new techniques that we can study and understand at a deeper level so as to mitigate, or at least foresee, some of these failure modes.
@@
Our work is not without limitations. In practice, estimating the conditional entropy $H(Y|X)$ is a difficult problem. That is part of the reason upper bounds might be useful, but that also means it is hard to evaluate how tight our upper bounds are. This also limits potential applications of our upper bounds like, for instance, estimating the expected prediction set size. Our experiments in Appendix G.3 rely on quantization to get a reasonable lower bound on the conditional entropy, and thus lower bound the expected prediction set size. Finally, in terms of our experimental results, we unfortunately have not been able to single out which of our new upper bounds performs best. The general trend we observe is that the DPI and model-based Fano bounds perform better on complex classification tasks, while simple Fano seems to excel in relatively simple tasks or where there is a higher risk of overfitting, like in the federated setting.
@@
In this paper, we focus our experiments on the classification setting, similarly to previous works on conformal training [8, 59]. However, in principle, our bounds pose no assumptions on the underlying prediction problem and we see the application of our results to the regression setting as a promise avenue for future work. In practice, only the simple Fano bound in (6) would not be directly applicable to the regression setting, but that is mainly because it assumes a uniform distribution over the output space and we get the, potentially infinite, $|\mathcal{Y}|$ term. The other two bounds we propose, DPI in (4) and model-based Fano in (5), do not include the $|\mathcal{Y}|$ and $|C(X)|$ terms and can be applied to regression problems as is.
@@ B.1 Exchangeability, Ranks and Quantiles
Formally, the concept of exchangeability can be defined as follows:
@@ C.1 Data processing inequalities for f-divergence
Theorem C.3 is the classical data processing inequality (DPI), which we restate below for the sake of completeness. The classical versions of DPI, stated in terms of mutual information, can be found in classical information theoretic text books like [13], while the generalization of DPI to $f$-divergence can be found in other works with a comprehensive survey in [49, 56].
@@
The proof can be found in standard textbooks in information theory see, for example, Chapter 7, Section 7.2. in [49] for a derivation of the DPI enunciated in the same form as above.
Next, we consider the application to conformal prediction.
@@
The proof simply follows from the data processing inequality by considering the random variable $E=\mathbf{1}(Y\in{\mathcal{C}}(x))$ and the conditional distribution $P_{E|X,Y}=\mathbb{E}_{{\mathcal{D}}_{cal}}[\mathbf{1}(Y\in{\mathcal{C}}(x))|X,Y]$. This construction can be further improved as follows.
@@
$D_{f}(P_{Y|X=x}||Q_{Y|X=x})$ $=D_{f}(P_{Y|X=x}P_{E|X=x,Y}||Q_{Y|X=x}P_{E|X=x,Y})$ (12)
$=D_{f}(P_{Y,E|X=x}||Q_{Y,E|X=x})$ (13)
and from the monotonicity property of the f-divergence [49] we have that
$D_{f}(P_{Y|X=x}||Q_{Y|X=x})$ $=D_{f}(P_{Y,E|X=x}||Q_{Y,E|X=x})$ (14)
$\geq D_{f}(P_{E|X=x}||Q_{E|X=x})$ (15)
$=d_{f}(P_{Y|X}(Y\in C(x)|X=x)||Q_{Y|X}(Y\in C(x)|X=x)).$ (16)
@@ C.4 Fano and Data Processing Inequalities for Conformal Prediction
First, from Fano’s inequality for list decoding, Theorem C.6, we get the next result ‘out-of-the-box’.
@@ C.5 Related Work on Information Theoretic Inequalities
The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s—see for instance [2, 14, 42]. A key $f$-divergence inequality is the data processing inequality, which was used in information theory to establish various upper bounds on the achievability of coding schemes for different tasks—see [56, 55] for an extensive survey. Moreover, the data processing inequality for $f$-divergences provides a unified way of obtaining many classical and new results, for example Fano’s inequality [18]. The tightness of data processing inequalities is discussed in terms of Bregman’s divergence in [12, 32] and in terms of $\chi^{2}$-divergence in [55].
When it comes to list decoding, there are a number of relevant inequalities in the literature. List decoding was introduced in context communication design by Elias [17] and Wozencraft [66]. For fixed list size, the information theoretic bounds on list decoding were obtained in [19] using error exponent analysis. A generalization of Fano’s inequality to list decoding was given in [15] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [51]. For fixed list size, stronger inequalities, some based on Arimoto-Rényi conditional entropy, were presented in [57]. Variable size list decoding was discussed in [55] using the notion of $E_{\gamma}$ resolvability first introduced in [50] related to the dependence testing bound. It was used again in the context of channel resolvability in [33], where some relevant inequalities have been obtained and discussed. A selection of the most relevant inequalities for list decoding can be found in [55].
@@ Appendix D Proofs of Main Theoretical Results
In this section, we provide the proofs to our main results, namely the DPI bound in Proposition 3.1, the model-based Fano bound in Proposition 3.2, and the simple Fano bound in Corollary 3.1. For notational convenience, we use the shorthand $\alpha_{n}=\alpha-\nicefrac{{1}}{{n+1}}$ in most of the steps of the derivations.
@@ Appendix F Conformal Training
Split conformal prediction (SCP) [45] has quickly become a popular framework for uncertainty estimation, largely thanks to its computational efficiency. One only needs access to a separate calibration data set to derive prediction sets with valid marginal coverage from any pretrained model. Given that training new machine learning models is becoming ever more time-consuming and expensive with new, larger architectures, this ability to apply conformal prediction to existing models is invaluable in a number of applications. Yet, it is reasonable to expect that the performance of the final set predictor could be improved if the conformal prediction process were to be accounted for during training of the model as well, steering the model towards better predictive efficiency, as it were. That is the motivation behind the line of work that we broadly refer to as conformal training [8, 59]. In a nutshell, conformal training introduces a differentiable, and hence approximate, conformal prediction step during training so that one can directly optimize for desired properties of the set predictor, most notably its predictive efficiency.
To the best of our knowledge, Bellotti [8] were the first to advance the idea of conformal training. The central step of their approach is to relax the prediction set defined by the model $f$ and define “soft” prediction sets $\hat{\mathcal{C}}_{f}(x)$, which contain each of the labels $y\in{\mathcal{Y}}$ with a certain probability. That is, if ${\mathcal{C}}_{f}(x,y)\in\{0,1\}$ is the hard assignment of label $y$ to ${\mathcal{C}}(x)$, a corresponding soft version of this assignment can be defined as
@@
Bellotti [8] then goes on to propose the following loss functions which are computed for each training batch ${\mathcal{B}}$
@@
Importantly, Bellotti [8] argues the choice of the threshold $\hat{q}$ in (29) is immaterial since the model can learn to shift its logit values accordingly to match the constraints in ${\mathcal{L}}_{coverage}(f)$. We can then directly optimize (30) via stochastic gradient descent methods during training since it is fully differentiable with respect to the model parameters.
@@
$\hat{\mathcal{C}}_{f}(x,y):=\sigma\left(\frac{\text{Quantile}(\alpha;{\mathcal{B}}_{cal})-s(x,y)}{T}\right).$
@@
This version of their approach, which we refer to as ConfTr, only optimizes the size loss, but Stutz et al. [59] also proposes another variant which includes a classification loss term as follows
${\mathcal{L}}_{class}(\hat{\mathcal{C}}_{f}(x),y)=\sum_{\hat{y}\in{\mathcal{Y}}}L_{y,\hat{y}}\left[\left(1-\hat{\mathcal{C}}_{f}(x,y)\right)\delta[\hat{y}=y]+\hat{\mathcal{C}}_{f}(x,y)\delta[\hat{y}\neq y]\right],$
@@
Our own approach to conformal training follows the same recipe from [59], i.e., we also simulate a split conformal prediction step during training by splitting each training batch into two and using differentiable sorting operators. The key difference is in how we define the training objectives, which we derive from first principles and standard information theory inequalities. Not only do our upper bounds, DPI (4), MB Fano (5) and Fano (6), outperform the ConfTr objectives in a number cases, but they also do away with a few hyperparameters. Namely, the function $g$ in the ${\mathcal{L}}_{size}$, and hyperparameters controlling the relative importance of ${\mathcal{L}}_{size}$ and ${\mathcal{L}}_{class}$.
@@ F.1 Deriving Conformal Training from Fano’s bound
Through Proposition 3.2, we can connect Fano’s bound for list decoding to the size loss from [59], proposed for conformal training. More specifically, by assuming a uniform distribution for $Q$ we can show that

### v3

@@ 7 Experiments
In this section, we empirically study two applications of our theoretical results, namely conformal prediction with side information and conformal training with our upper bounds on the conditional entropy as optimization objectives. We focus our experiments on classification tasks since this is the most common setting in previous works in conformal training [13, 8, 61].
@@ 7.1 Conformal Training
In Table 1, we report the empirical inefficiency on test data considering two SCP methods, threshold CP with probabilities (or THR) [56] and APS [55]—see Appendix G.1.1 for results with RAPS [4]. In all cases, our upper bounds proved effective loss functions to train efficient classifiers end-to-end and from scratch. For the simpler data sets (MNIST, Fashion-MNIST and EMNIST), all conformal training methods achieved similar results, but both ConfTr methods proved less consistent. This is noticeable in the oftentimes sharp difference in performance between THR and APS, since even after fine-tuning the hyperparameters (see Appendix G) some of the models failed to converge properly. For the remaining and more challenging data sets, both ConfTr variants lagged behind, probably because they do not provide a strong enough signal to train ResNets from scratch (on CIFAR data sets, Stutz et al. [61] only used ConfTr to fine tune pretrained models). A similar observation applies to the simple Fano bound (6), where the relaxed prediction set size is the only learning signal.
Method $\alpha=0.01$ $\alpha=0.05$ $\alpha=0.1$
THR APS THR APS THR APS
[…]
@@ 7.3 Federated Learning (FL)
Method EMNIST CIFAR 10 CIFAR 100
APS APS${}_{+\text{SI}}$ APS APS${}_{+\text{SI}}$ APS APS${}_{+\text{SI}}$
[…]
@@
One marked difference between Tables 1 and 4 is that the simple Fano bound (6), which lagged behind the DPI bound and its model-based counterpart in the centralized setting, achieved the best results on the federated setting. We hypothesize this could be due to overfitting of the local optimization procedures to the individual data distribution of each device, which hurts the convergence of the global model. This effect is exacerbated on CIFAR100, where we have 500 devices, each of which with very few data points. The simple Fano bound is less vulnerable to such overfitting since it relies on the main classifier to a much lesser degree. Finally, in almost all cases, our bounds outperformed the CE loss, reassuring the potential of conformal training. Moreover, the inclusion of side information reduced inefficiency in all settings, and markedly so in a few instances. This confirms the effectiveness of our side information approach in a complex and practically relevant scenario, like federated learning.
@@ 8 Conclusion
In this work, we established a link between notions of uncertainty coming from conformal prediction and information theory (or more precisely variable-size list decoding). We proved that one can use split conformal prediction methods to upper bound the conditional entropy of the target variable given the inputs, and that these upper bounds form principled objectives for conformal training. We empirically validated our approach to conformal training, with strong results in both centralized and federated settings. Furthermore, the information-theoretic perspective also offers a simple yet rigorous approach to incorporate side information into conformal prediction, which we experimentally show leads to better predictive efficiency. To the best of our knowledge, this is the first attempt at connecting information theory and conformal prediction. Given the limited communication between these two research communities thus far, we expect our work to incite a fruitful exchange of not only ideas but also theory and algorithms between these two research domains. In this paper, we concentrated our exposition and experiments in classification tasks, but we see an extension of our methods to the regression setting, as a particularly promising avenue for future work.
@@ Appendix A Broader Impact, Limitations and Other Remarks
This work explores the connection between conformal prediction and information theory, with the end goal of advancing the state of the art of uncertainty quantification in machine learning. With that in mind, we believe its potential societal consequences are chiefly positive, since our work might contribute to a larger adoption of uncertainty estimates, in a number of safety-critical applications. Notwithstanding, conformal prediction, like any other uncertainty estimation method, should be applied with care and a proper understanding of the provided guarantees so as not to create an illusion of safety when there is none. That is why our work aims to develop new methods and algorithms from first principles, with the hope of providing new techniques that we can study and understand at a deeper level so as to mitigate, or at least foresee, some of their failure modes.
@@
Our work is not without limitations. In practice, estimating the conditional entropy $H(Y|X)$ is a difficult problem. That is part of the reason upper bounds might be useful, but that also means it is hard to evaluate how tight our upper bounds are. This also limits potential applications of our upper bounds like, for instance, estimating the expected prediction set size. Our experiments in Appendix G.3 rely on quantization to get a reasonable lower bound on the conditional entropy, and thus lower bound the expected prediction set size. Finally, in terms of our experimental results, we unfortunately have not been able to single out which of our new upper bounds performs best. The general trend we observe is that the DPI and model-based Fano bounds perform better on complex classification tasks, while simple Fano seems to excel in relatively simple tasks or where there is a high risk of overfitting, like in the federated setting.
@@
In this paper, we focus our experiments on the classification setting, similarly to previous works on conformal training [13, 8, 61]. However, in principle, our bounds pose no assumptions on the underlying prediction problem and we see the application of our results to the regression setting as a promising avenue for future work. In practice, only the simple Fano bound in (6) would not be directly applicable to the regression setting, but that is mainly because it assumes a uniform distribution over the output space and we get the, potentially infinite, $|\mathcal{Y}|$ term. The other two bounds we propose, DPI in (4) and model-based Fano in (5), do not include the $|\mathcal{Y}|$ and $|C(X)|$ terms and can be applied to regression problems as is.
@@ B.1 Exchangeability, Ranks and Quantiles
Formally, the concept of exchangeability can be defined as follows.
@@ C.1 Data processing inequalities for f-divergence
Theorem C.3 is the classical data processing inequality (DPI), which we restate below for the sake of completeness. The classical versions of DPI, stated in terms of mutual information, can be found in classical information theoretic text books like [14], while the generalization of DPI to $f$-divergences can be found in other works with a comprehensive survey in [51, 58].
@@
The proof can be found in standard textbooks in information theory see, for example, Chapter 7, Section 7.2 in [51] for a derivation of the DPI enunciated in the same form as above. Next, we consider the application of the DPI to conformal prediction.
@@
Proof.
Consider the random variable $E=\mathbf{1}(Y\in{\mathcal{C}}(x))$ denoting the event of valid coverage, and the conditional distribution $P_{E|X,Y}=\mathbb{E}_{{\mathcal{D}}_{cal}}[\mathbf{1}(Y\in{\mathcal{C}}(x))|X,Y]$. Note that $P_{E|X,Y}$ maps distributions $P_{X,Y}$ and $Q_{X,Y}$ to $P_{E}$ and $Q_{E}$, resp. From here, the result follows directly from Theorem C.3. ∎
The construction in Theorem C.4 can be further improved as follows.
@@
$D_{f}(P_{Y|X=x}||Q_{Y|X=x})$ $=D_{f}(P_{Y|X=x}P_{E|X=x,Y}||Q_{Y|X=x}P_{E|X=x,Y})$
$=D_{f}(P_{Y,E|X=x}||Q_{Y,E|X=x})$
and from the monotonicity property of f-divergences [51] we have that
$D_{f}(P_{Y|X=x}||Q_{Y|X=x})$ $=D_{f}(P_{Y,E|X=x}||Q_{Y,E|X=x})$
$\geq D_{f}(P_{E|X=x}||Q_{E|X=x})$
$=d_{f}(P_{Y|X}(Y\in C(x)|X=x)||Q_{Y|X}(Y\in C(x)|X=x)).$
@@ C.4 Fano and Data Processing Inequalities for Conformal Prediction
First, from Fano’s inequality for list decoding, Theorem C.6, we get the next result “out-of-the-box”.
@@ C.5 Related Work on Information Theoretic Inequalities
The use of $f$-divergences goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s—see for instance [2, 15, 44]. A key $f$-divergence inequality is the data processing inequality, which was used in information theory to establish various upper bounds on the achievability of coding schemes for different tasks—see [58, 57] for an extensive survey. Moreover, the data processing inequality for $f$-divergences provides a unified way of obtaining many classical and new results, for example Fano’s inequality [19]. The tightness of data processing inequalities is discussed in terms of Bregman’s divergence in [12, 34] and in terms of $\chi^{2}$-divergence in [57].
When it comes to list decoding, there are a number of relevant inequalities in the literature. List decoding was introduced in the context of communication design by Elias [18] and Wozencraft [68]. See also [22] for a more recent overview of list decoding. For fixed list size, the information theoretic bounds on list decoding were obtained in [20] using error exponent analysis. A generalization of Fano’s inequality to list decoding was given in [16] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [53]. For fixed list size, stronger inequalities, some based on Arimoto-Rényi conditional entropy, were presented in [59]. Variable size list decoding was discussed in [57] using the notion of $E_{\gamma}$ resolvability first introduced in [52] related to the dependence testing bound. It was used again in the context of channel resolvability in [35], where some relevant inequalities have been obtained and discussed. A selection of the most relevant inequalities for list decoding can be found in [57].
@@ Appendix D Proofs of Main Theoretical Results
In this section, we provide the proofs of our main results, namely the DPI bound in Proposition 3.1, the model-based Fano bound in Proposition 3.2, and the simple Fano bound in Corollary 3.1. For notational convenience, we use the shorthand $\alpha_{n}=\alpha-\nicefrac{{1}}{{n+1}}$ in most of the steps of the derivations.
@@ Appendix F Conformal Training
Split conformal prediction (SCP) [47] has quickly become a popular framework for uncertainty quantification, largely thanks to its computational efficiency. One only needs access to a separate calibration data set to derive prediction sets with valid marginal coverage from any pretrained model. Given that training new machine learning models is becoming ever more time-consuming and expensive with new, larger architectures, this ability to apply conformal prediction to existing models is invaluable in a number of applications. Yet, it is reasonable to expect that the performance of the final set predictor could be improved if the conformal prediction process were to be accounted for during training of the model as well, steering the model towards better predictive efficiency, as it were. That is the motivation behind the line of work that we broadly refer to as conformal training [13, 8, 61]. In a nutshell, conformal training introduces a differentiable, and hence approximate, conformal prediction step during training so that one can directly optimize for desired properties of the set predictor, most notably its predictive efficiency. In what follows we give an overview of conformal training, focusing on classification tasks.
The idea of conformal training has been proposed concomitantly in [8, 13]. Here we follow the approach in [8], where the key idea is to relax the prediction set defined by the model $f$ and define “soft” prediction sets $\hat{\mathcal{C}}_{f}(x)$, which contain each of the labels $y\in{\mathcal{Y}}$ with a certain probability. That is, if ${\mathcal{C}}_{f}(x,y)\in\{0,1\}$ is the hard assignment of label $y$ to ${\mathcal{C}}(x)$, a corresponding soft version of this assignment can be defined as
@@
Bellotti [8] then proposes the following loss functions which are computed for each training batch ${\mathcal{B}}$
@@
Importantly, Bellotti [8] argues the choice of the threshold $\hat{q}$ in (24) is immaterial since the model can learn to shift its outputs (in logit space) accordingly to match the constraints in ${\mathcal{L}}_{coverage}(f)$. We can then directly optimize (25) via stochastic gradient descent methods during training since it is fully differentiable with respect to the model parameters.
@@
$\hat{q}=\text{Quantile}(1-\alpha;\{s_{f}(x,y):(x,y)\in{\mathcal{B}}_{cal}\})$
@@
This version of their approach, which we refer to as ConfTr, only optimizes the size loss, but Stutz et al. [61] also proposed another variant which includes a classification loss term as follows
${\mathcal{L}}_{class}(f)=\frac{1}{|{\mathcal{B}}|}\sum_{(x,y)\in{\mathcal{B}}}\sum_{\hat{y}\in{\mathcal{Y}}}L_{y,\hat{y}}\left[\left(1-\hat{\mathcal{C}}_{f}(x,y)\right)\delta[\hat{y}=y]+\hat{\mathcal{C}}_{f}(x,y)\delta[\hat{y}\neq y]\right],$ (26)
@@
Our own approach to conformal training follows the same recipe from [61], i.e., we also simulate a split conformal prediction step during training by splitting each training batch into two and using differentiable sorting operators (see Algorithm 1). The key difference is in how we define the training objectives, which we derive from first principles and standard information theory inequalities. Not only do our upper bounds, DPI (4), MB Fano (5) and Fano (6), outperform the ConfTr objectives in many cases, but they also do away with a few hyperparameters. Namely, the function $g$ in ${\mathcal{L}}_{size}$, and hyperparameters controlling the relative importance of ${\mathcal{L}}_{size}$ and ${\mathcal{L}}_{class}$.
@@ F.1 Deriving Conformal Training from Fano’s bound
Through Proposition 3.2, we can connect Fano’s bound for list decoding to the size loss from [61], proposed for conformal training. Assuming a uniform distribution for $Q$ we can show that